OpenStudy (anonymous):

PLEASE HELP h= -16t^2 +44t + 5 How high is the ball after 2 seconds?

OpenStudy (texaschic101):

h = -16(2^2) + 44(2) + 5 h = -16(4) + 88 + 5 h = 64 + 88 + 5 h = 157